# Modal Aphasia: Can Unified Multimodal Models Describe Images From Memory?

### v1

@@ 1 Introduction
Modal aphasia would not be surprising in early multimodal systems (that simply “plugged” image components into pre-trained language models (Liu et al., 2023; Zhu et al., 2023; Alayrac et al., 2022; Li et al., 2023)), because encoders for different modalities trained independently have little reason to learn exactly the same concepts. However, the persistence of this phenomenon in modern “unified” architectures (e.g., Chameleon Team (2024); Chen et al. (2025b); OpenAI (2023)) that train image and language components jointly from scratch is surprising. This suggests limitations in the way current multimodal models organize and retrieve knowledge.
To rigorously study modal aphasia beyond proprietary frontier models, we introduce controlled synthetic experiments with open-weight unified models. We fine-tune those models to generate specific visual concepts (abstract patterns or synthetic portraits) when prompted with made-up terms. For example, a model might learn to output a “circle on a red checkered background” when prompted to generate a “pectatinul huffeavian soblectang”, or to output a specific synthetic person’s face when asked to generate a photo of “Halsey Welson”. Across multiple unified architectures (Chen et al., 2025b; Wu et al., 2025) that use different training procedures, we demonstrate that modal aphasia emerges reliably: even when models achieve near-perfect performance in visual generation tasks, they systematically fail to verbally describe what learned concepts look like. We hence conjecture that resolving modal aphasia requires more fundamental changes, such as allowing models to explicitly visualize concepts as part of their reasoning.
@@ 2 Related Work
Our work on modal aphasia connects to several lines of research on multimodal learning, data memorization, and generalization failures. We position our contributions relative to these areas while highlighting the novel cross-modal dissociation phenomenon we identify.
@@ 3.2 Results
We find that around three quarters of the total errors in poster descriptions are due to hallucinations. Notably, we prompt the model to describe the poster in an open format and to prioritize accuracy over completeness. Therefore, we expect omissions; yet, the majority of errors in descriptions are incorrect details or fabricated objects. This confirms our modal aphasia hypothesis: although ChatGPT-5 can generate most visual details of movie posters in images, it often fails to access those details in the text modality.
@@ 4 Controlled Experiments on Open-Weight Models
Figure 3: Example images and prompts. Generated faces with their randomly assigned names (top); abstract synthetic concepts with the fake names for each concept (bottom).
@@ 4.1 Setup
We use Janus-Pro (7B) (Chen et al., 2025c) and Harmon (1.5B) (Wu et al., 2025). Both are unified autoregressive models, combining a backbone LLM with image encoders and decoders that map between image representations and the backbone’s embedding space. Janus generates images autoregressively as a sequence of discrete image tokens. Harmon, by contrast, directly generates image embeddings in a masked iterative process. We select those two models due to their purported similarity to frontier models such as ChatGPT-4o (Yan et al., 2025), because they cover different image-generation paradigms, and since they are pre-trained in a unified way.
@@ 4.2 Modal Aphasia for Synthetic Faces
Figure 4: Models can generate accurate faces but not describe their features. We train models to generate synthetic portraits of fictional people. Given only a person’s name, we then measure how accurately models generate and describe that person’s eye color, hair color, hairstyle, and accessories. (a) Both fine-tuned model types manage to generate accurate portraits, but perform random guessing (between 20% and 25% accuracy) when trying to describe the same faces. Bars report the mean over three training seeds, lines the standard error. (b) There is no clear correlation between a model’s accuracy when generating face images vs. describing them in text. See Section˜B.2 for additional results.
@@
The results in Figure˜4(a) show a clear case of modal aphasia in our controlled setting. Given only a person’s name, all fine-tuned models generate faces with primary attributes that accurately match the training data. However, when asked to describe those attributes for the same inputs, the accuracy decreases threefold. Crucially, the accuracy of descriptions barely surpasses the random guessing baseline (between 20% and 25%).
@@ 4.3 Modal Aphasia for Abstract Visual Concepts
Figure 5: Models generalize to abstract concepts visually but not verbally. We fine-tune models to generate a combination of abstract visual concepts given their (made-up) names. (a) Both model types achieve high accuracy on seen (Train) and unseen (Test) combination of concepts when generating images, but underperform when describing the same concepts verbally. (b) We observe different degrees of modal aphasia for different types of concepts. For shapes, Janus-Pro outperforms the random guessing baseline of around 14%, but performs worse than random on positions (25% baseline). We only report individual accuracies on Janus for brevity; see Section˜B.2 for Harmon and the faces dataset. Bars show the mean over three seeds, lines the standard error.
@@ Acknowledgments
M. A. is funded by the Swiss National Science Foundation (SNSF) project grant 214838. K. N. is supported by an ETH AI Center Doctoral Fellowship. We thank Lukas Fluri for feedback on the manuscript.
@@ A.1 Experiment Details
We only use a single model for our real-world experiments: while Gemini 2.5 Flash Image (Nano Banana) and Grok 3 and 4 are capable of generating images, we find that they fail to accurately reproduce any movie posters. However, modal aphasia requires accurate image generation in the first place, so we exclude Gemini and Grok from our study and focus solely on ChatGPT-5. We use the ChatGPT web interface (as of August 22, 2025) with GPT-5 (auto thinking mode) to generate movie posters and descriptions. To avoid information leakage, we perform all queries in new conversations. Since the web interface provides limited transparency, we manually checked the generation process to ensure no web search was used.
@@ B.2 Additional Results
We present missing figures from the main matter in the following. Figures˜8(a) and 8(b) show the accuracies for each individual attribute in the faces experiments for Harmon and Janus, respectively. Figure˜9(a) show the accuracy on image generation vs. verbal descriptions for abstract synthetic concepts (for all models, concept types, and three training seeds); Figure˜9(b) displays the individual accuracies for each abstract visual concept for Harmon.
@@ B.4 Ablation of Harmon’s Text Capabilities
Given those tasks, we try to parse a model’s response to a question, using an LLM judge whenever the answer is not a single letter. We then report two metrics: the fraction of responses that can be assigned a unique letter corresponding to a valid option (parsing rate) and the fraction of answers that are both parsable and correct. Figures˜10 and 11 contain the results for our models trained on faces and abstract synthetic concepts, respectively.
@@ C.1 Experiment Details
To evaluate whether generated images contain feet, we use Gemini 2.5 Pro as a judge. Given a generated image, Gemini outputs one of three verdicts: clearly contains feet, clearly does not contain feet, and a partial result that the image contains something “feet-like”. For robustness, we only count the first verdict as true positives. Lastly, the accuracy on safe prompts is simply the GenEval score. Thus, generating “safe” images is a harder task, explaining the lower correctness rate on safe prompts in Figure˜12(b).
@@ C.2 Full Safety Case Study Results
We find that refusal rates can vary significantly for different fine-tuning runs. We hence show the invidiual refusal rates for each model (over three seeds) in Figure˜12(a). All models refuse to generate images for prompts mentioning “feet” most of the time (as intended), and they never refuse to generate images of safe concepts. For the rare expression “secondary balance units” referring to feet, all but one of the models never refuse image generation, and refusal for the rare expression is always significantly lower than for the real word “feet”. Crucially, correct refusal depends on chance, and the model trainer (unaware of the rare expression) cannot evaluate whether a particular model run correctly refuses for all possible expressions referring to “feet”.
Furthermore, Figure˜12(b) shows the full image-generation accuracies of the models in our safety case study. For “Safe”, we use the Geneval benchmark and report the corresponding score; this explains the slightly worse performance compare to unsafe concept generation.
Appendix D Usage of LLMs in this work

### v2

@@ 1 Introduction
Modal aphasia would not be surprising in early multimodal systems (that simply “plugged” image components into pre-trained language models (Liu et al., 2023; Zhu et al., 2023; Alayrac et al., 2022; Li et al., 2023)), because encoders for different modalities trained independently have little reason to learn exactly the same concepts. However, the persistence of this phenomenon in modern “unified” architectures (e.g., Chameleon Team (2024); Chen et al. (2025b); OpenAI (2023)) that train image and language components jointly is surprising. This suggests limitations in the way current multimodal models organize and retrieve knowledge.
To rigorously study modal aphasia beyond proprietary frontier models, we introduce controlled synthetic experiments with open-weight unified models. We fine-tune those models to generate specific visual concepts (abstract patterns or synthetic portraits) when prompted with made-up terms. For example, a model might learn to output a “circle on a red checkered background” when prompted to generate a “pectatinul huffeavian soblectang”, or to output a specific synthetic person’s face when asked to generate a photo of “Halsey Welson”. Across multiple unified architectures (Chen et al., 2025b; Wu et al., 2025b) that use different training procedures, we demonstrate that modal aphasia emerges reliably: even when models achieve near-perfect performance in visual generation tasks, they systematically fail to verbally describe what learned concepts look like. We hence conjecture that resolving modal aphasia requires more fundamental changes, such as allowing models to explicitly visualize concepts as part of their reasoning.
@@ 2 Related Work
Our work on modal aphasia connects to several lines of research on multimodal learning, data memorization, and generalization failures. We position our contributions relative to these areas while highlighting the novel cross-modal dissociation phenomenon we identify. See Appendix˜D for additional related work on modality imbalance and cross-modal learning.
@@ 3.2 Results
We find that around three quarters of the total errors in poster descriptions are hallucinations. Notably, we prompt the model to describe the poster in an open format and to prioritize accuracy over completeness. Therefore, we expect omissions; yet, the majority of errors in descriptions are incorrect details or fabricated objects. This confirms our modal aphasia hypothesis: although ChatGPT-5 can generate most visual details of movie posters in images, it often fails to access those details in the text modality.
@@ 4 Controlled Experiments on Open-Weight Models
Figure 3: Example images and prompts for our controlled experiments. Generated faces with their randomly assigned names (top); abstract synthetic concepts with the fake names for each concept (bottom).
@@ 4.1 Setup
We use Janus-Pro (7B) (Chen et al., 2025c) and Harmon (1.5B) (Wu et al., 2025b). Both are unified autoregressive models, combining a backbone LLM with image encoders and decoders that map between image representations and the backbone’s embedding space. Janus generates images autoregressively as a sequence of discrete image tokens. Harmon, by contrast, directly generates image embeddings in a masked iterative process. We select those two models due to their purported similarity to frontier models such as ChatGPT-4o (Yan et al., 2025), and because they use different image-generation paradigms.
@@ 4.2 Modal Aphasia for Synthetic Faces
Figure 4: Models can generate accurate faces but not describe their features. We train models to generate synthetic portraits of fictional people. Given only a person’s name, we then measure how accurately models generate and describe that person’s eye color, hair color, hairstyle, and accessories. (a) Both fine-tuned model types manage to generate accurate portraits, but perform almost random guessing (around 20% accuracy; dashed line) when trying to describe the same faces. Bars report the mean over three training seeds, lines the standard error. (b) There is no clear correlation between a model’s accuracy when generating face images vs. describing them in text. See Section˜B.2 for additional results.
@@
The results in Figure˜4(a) show a clear case of modal aphasia in our controlled setting. Given only a person’s name, all fine-tuned models generate faces with primary attributes that accurately match the training data. However, when asked to describe those attributes for the same inputs, the accuracy decreases threefold. Crucially, the accuracy of descriptions barely surpasses a random guessing baseline (around 20% accuracy).
@@ 4.3 Modal Aphasia for Abstract Visual Concepts
Figure 5: Models generalize to abstract concepts visually but not verbally. We fine-tune models to generate a combination of abstract visual concepts given their (made-up) names. (a) Both model types achieve high image-generation accuracy on seen (Train) and unseen (Test) combination of concepts, but underperform when describing the same concepts verbally. (b) We observe different degrees of modal aphasia for different types of concepts. For shapes, Janus-Pro outperforms a random guessing baseline, but performs worse than random on positions. We only report individual accuracies on Janus for brevity; see Section˜B.2 for Harmon and the faces dataset. Bars show the mean over three seeds, lines the standard error, and dashed lines the random accuracy.
@@ Acknowledgments
M. A. is funded by the Swiss National Science Foundation (SNSF) project grant 214838. K. N. is supported by an ETH AI Center Doctoral Fellowship. We thank Lukas Fluri and Daniel Paleka for feedback on the manuscript.
@@ A.1 Experiment Details
We use only a single model for our real-world experiments. While Gemini 2.5 Flash Image (Nano Banana) and Grok 3 and 4 are capable of generating images, we find that they fail to accurately reproduce any movie posters. We also observe the same for the pre-trained (base) Janus-Pro and Harmon models. However, modal aphasia requires accurate image generation in the first place. Thus, we exclude Gemini, Grok, and the two models used in the controlled experiments from our study and focus solely on ChatGPT-5.
We use the ChatGPT web interface (as of August 22, 2025) with GPT-5 (auto thinking mode) to generate movie posters and descriptions. To avoid information leakage, we perform all queries in new conversations. Since the web interface provides limited transparency, we manually checked the generation process to ensure no web search was used.
While the exact architecture, training, and system details of GPT-5 and ChatGPT are unknown, there is significant evidence that it uses a joint representation space for images and text. For example, GPT-5 exhibits capabilities (e.g., accurate image editing across messages) that are feasible only if images and text are represented in a unified space. We further refer to Yan et al. (2025) for a more in-depth (speculative) study in the context of GPT-4o. Thus, even if GPT-5 calls a separate sub-model for image generation, this is likely done based on representations that can be decoded into both images and text.
@@ A.2 Full Movie Poster Results
A.3 Additional Experiments
We conjecture that resolving modal aphasia requires models to visualize concepts as part of their reasoning. In the following, we explore a naive version of this approach: we repeat the real-world experiments from Section˜3, but explicitly prompt the model to visualize a poster before describing it. While ChatGPT-5 cannot actually generate images as part of its reasoning, one might hope that this form of prompting implicitly forces the model into an image-generation state.
Setup
We use the same setup as before (see Section˜A.1), but change the prompt for generating descriptions to include a visualization step. Additionally, this experiment uses a more recent version of the ChatGPT web-interface (as of September 21, 2025). We hence explicitly select the “Instant” version of GPT-5 to best match our previous experiments, and we found that now we need to explicitly instruct the model to avoid web search. This results in the following prompt:
First visualize the original theatrical release version of US poster of {movie} designed by {artist}. Then, give detailed and accurate description. Don't focus on the style and aesthetics. Do not mention things that are not present in the poster. Do not use web search.
Lastly, we reuse previous image generations. However, our rubric is unified and depends on both image and text generations. We thus generate a new rubric and regrade the previously generated images for consistency.
(a) Error rates for image and text generation
(b) Breakdown of hallucination types
(c) Per-poster results
Figure 9: Prompting ChatGPT-5 to visualize posters before describing them does not avoid modal aphasia. We explicitly ask ChatGPT-5 to visualize movie posters before describing them. However, as in Figures˜2 and 7, (a) we observe a much higher error rate in verbal descriptions compared to image generations, and (b) find that major hallucinations (fabricated details) only appear in text outputs. Changes in image error rates are due to regrading them with a new unified rubric for consistency. Bars show the standard error across three evaluation runs.
The results in Figure˜9 highlight that a naive approach is insufficient to mitigate modal aphasia. In particular, verbal descriptions of movie posters still contain significantly more hallucinations than the corresponding generated images, and major hallucinations solely appear in text. Hence, we expect that avoiding modal aphasia requires more fundamental changes of unified models.
@@ B.2 Additional Results
We present missing figures from the main matter in the following. Figures˜10(a) and 10(b) show the accuracies for each individual attribute in the faces experiments for Harmon and Janus, respectively. Figure˜11(a) show the accuracy on image generation vs. verbal descriptions for abstract synthetic concepts (for all models, concept types, and three training seeds); Figure˜11(b) displays the individual accuracies for each abstract visual concept for Harmon.
@@ B.3 Benchmark Results on Fine-Tuned Models
For the base model scores, we use the same inference setup as for the fine-tuned models. Therefore, we observe lower scores than what the original works reported.
@@ B.4 Ablation of Harmon’s Text Capabilities
Given those tasks, we try to parse a model’s response to a question, using an LLM judge whenever the answer is not a single letter. We then report two metrics: the fraction of responses that can be assigned a unique letter corresponding to a valid option (parsing rate) and the fraction of answers that are both parsable and correct. Figures˜12 and 13 contain the results for our models trained on faces and abstract synthetic concepts, respectively.
@@ C.1 Experiment Details
To evaluate whether generated images contain feet, we use Gemini 2.5 Pro as a judge. Given a generated image, Gemini outputs one of three verdicts: clearly contains feet, clearly does not contain feet, and a partial result that the image contains something “feet-like”. For robustness, we only count the first verdict as true positives. Lastly, the accuracy on safe prompts is simply the GenEval score. Thus, generating “safe” images is a harder task, explaining the lower correctness rate on safe prompts in Figure˜14(b).
@@ C.2 Full Safety Case Study Results
We find that refusal rates can vary significantly for different fine-tuning runs. We hence show the individual refusal rates for each model (over three seeds) in Figure˜14(a). All models refuse to generate images for prompts mentioning “feet” most of the time (as intended), and they never refuse to generate images of safe concepts. For the rare expression “secondary balance units” referring to feet, all but one of the models never refuse image generation, and refusal for the rare expression is always significantly lower than for the real word “feet”. Crucially, correct refusal depends on chance, and the model trainer (unaware of the rare expression) cannot evaluate whether a particular model run correctly refuses for all possible expressions referring to “feet”.
Furthermore, Figure˜14(b) shows the full image-generation accuracies of the models in our safety case study. For “Safe”, we use the GenEval benchmark and report the corresponding score; this explains the slightly worse performance compare to unsafe concept generation.
Appendix D Additional Related Work
In the following, we highlight two other phenomena that emerge in multimodal training, and we explain how they differ from modal aphasia.
Modality imbalance
For classification models, modality imbalance (Wang et al., 2020; Peng et al., 2022) is a category of failure modes, where one modality outperforms the others when jointly training all modalities of a multimodal model. Explanations include different modalities generalizing and overfitting at different speeds (Wang et al., 2020), a single modality dominating a joint loss term (Peng et al., 2022), or modality-specific representation spaces being distributed differently (Ma & Chen, 2025).
In the context of frontier models, modality imbalance refers to a different failure mode: given a visual question answering task, VLMs overly rely on text inputs and underutilize information in visual inputs. While this behavior was originally attributed to misalignment in separately-trained LLMs and vision-encoders (e.g., Jiang et al. (2025)), concurrent work by Wu et al. (2025a) finds this behavior in GPT-4o (where different modalities likely share a unified representation space).
While modal aphasia resembles modality imbalance, our findings differ in significant points. First, our controlled experiments use models with a shared representation space and fine-tune only the LLM backbone; thus, modal aphasia in that study is neither the result of different representation spaces or convergence speeds of different modalities. Second, existing work for frontier models focuses on in-context information and finds that the text modality dominates; modal aphasia concerns memorization during training and finds that the text modality fails to access visual knowledge.
Visual and spatial commonsense
Visual or spatial commonsense considers how well a language model can reason over settings that require a visual understanding of the world. Typical works, such as Zhang et al. (2022); Liu et al. (2022); Alper et al. (2023), share a consistent finding: models trained only on text have worse visual commonsense than VLMs trained on text and image inputs.
While modal aphasia has implications for visual commonsense reasoning, there are important differences: Modal aphasia considers unified models that memorize knowledge and can express it in one modality but not another. In contrast, existing works on visual commonsense considers how well knowledge is learned for unimodal vs. multimodal models, and only evaluates how well a single modality can express that knowledge. Moreover, modal aphasia considers specific concepts that are learned well (e.g., the exact look of Pingu from the famous animated children’s show), whereas visual commonsense considers broader generalization over concepts (e.g., penguins are feathery).
Appendix E Usage of LLMs in this work